OpenStudy (anonymous):

What is the simplified form of the reciprocal of: squareroot 2 + squareroot 18?

OpenStudy (anonymous):

u would have 1 on top of both fractions.then u simplify square root 20 to 2 square root 5 after adding like terms. you end up with ::: 1 over 2 square root of 5

OpenStudy (anonymous):

thats not one of the multiple choice options

OpenStudy (anonymous):

ok so how is it written exactly?

myininaya (myininaya):

\[\sqrt{2}+\sqrt{18}=\sqrt{2}+\sqrt{9}\sqrt{2}=\sqrt{2}+3\sqrt{2}=(1+3)\sqrt{2}=4\sqrt{2}\]

myininaya (myininaya):

oops missed the reciprocal part

myininaya (myininaya):

\[\frac{1}{4\sqrt{2}}\]

myininaya (myininaya):

rationalize the denominator and you will have your answer

OpenStudy (anonymous):

ok i see what you did. that actually helps me on my hw. so before you can add the like terms u have to break them down?

myininaya (myininaya):

\[\frac{1}{4\sqrt{2}}*\frac{2}{2}=\frac{\sqrt{2}}{4*2}=\frac{\sqrt{2}}{8}\]

myininaya (myininaya):

that 2/2 is sqrt2/sqrt2

myininaya (myininaya):

got it sparky?

OpenStudy (anonymous):

yah thanks guys